OpenStudy (irish_waffl3):

factor -6^4y^5-15x^3y^2+9x^2y^3

OpenStudy (irish_waffl3):

\[-6^4y^5-15x^3y^2+9x^2y^3\]

OpenStudy (irish_waffl3):

Idk how to medal, but I will if you tell me how

OpenStudy (welshfella):

look for the GCF (greatest common factor) for the 3 terms

OpenStudy (cas_fangirl_14):

Do you want step by step instructions or just the answer?

OpenStudy (irish_waffl3):

both

OpenStudy (welshfella):

first look ate the 3 numbers -6 , -15 and 9 - what number will divide into each of theses?

OpenStudy (irish_waffl3):

3

OpenStudy (welshfella):

No direct answers please

OpenStudy (irish_waffl3):

three is the number that will divide evenly into the three numbers.

OpenStudy (welshfella):

yes now look at the x terms

OpenStudy (irish_waffl3):

-15x, and 9x?

OpenStudy (welshfella):

is there an 'x' in each term?

OpenStudy (irish_waffl3):

yes all three, nevermind

OpenStudy (irish_waffl3):

6x, 9x, and -15x

OpenStudy (welshfella):

is there a mistake in the first term should it be 6x^4y^5 ?

OpenStudy (irish_waffl3):

yes

OpenStudy (irish_waffl3):

sorry

OpenStudy (irish_waffl3):

\[-6x^4y^5-15x^3y^2+9x^2y^3\]

OpenStudy (welshfella):

OK so we need the GCF of x^4 , x^3 and x^2

OpenStudy (irish_waffl3):

how do I find that?

OpenStudy (welshfella):

look at the lowest, x^2 - will it divide into x^3 and x^4?

OpenStudy (irish_waffl3):

yes

OpenStudy (irish_waffl3):

i think

OpenStudy (welshfella):

so thats your GCF now what about the terms in y?

OpenStudy (irish_waffl3):

how do I divide the x terms

OpenStudy (irish_waffl3):

im sorry, im lost

OpenStudy (irish_waffl3):

in alg

OpenStudy (irish_waffl3):

good help though

OpenStudy (welshfella):

x^3 / x^2 = x^(3-2) = x x^4 / x^2 = x^(4-2) = x^2 - one of the laws of indices.

OpenStudy (irish_waffl3):

ok, how long does it usually take to solve?

OpenStudy (welshfella):

not long once you know the basic rules

OpenStudy (irish_waffl3):

ok

OpenStudy (welshfella):

so for y^5 , y^2 and y^3 the GCF is y^2 so GCF of the whole expression is 3x^2y^2 now you divide this into the 3 terms - the result goes into parentheses

OpenStudy (irish_waffl3):

ok, 3x^2y^2

OpenStudy (irish_waffl3):

why 3x

OpenStudy (welshfella):

3x^2 - because 3 is gcf of numbers and x^2 is gcf of the x terms

OpenStudy (irish_waffl3):

ok

OpenStudy (welshfella):

and y6 is gcf for y terms

OpenStudy (irish_waffl3):

ok

OpenStudy (irish_waffl3):

so it would be 3x^2y^6 ?

OpenStudy (welshfella):

so first divide -6x^4y^5 by 3x^2y^2 we get -2x^2y^3

OpenStudy (irish_waffl3):

ok

OpenStudy (welshfella):

no its y^2 because thats the gcf of the y terms

OpenStudy (welshfella):

so we have 3x^2y^2 ( -2 x^2y^3 ) and there are 2 more terms to go into the parentheses

OpenStudy (irish_waffl3):

ok

OpenStudy (welshfella):

the second one is -15 x^3 y^2 / 3 x^2 y^2 what does that come to . Try to do the division.

OpenStudy (irish_waffl3):

ok

OpenStudy (welshfella):

first part is -15 ? 3 = -5

OpenStudy (welshfella):

* -15 / 3 = -5

OpenStudy (irish_waffl3):

so-5

OpenStudy (irish_waffl3):

sorry, i went to get a pen

OpenStudy (welshfella):

right now x^3 / x^2 ?

OpenStudy (irish_waffl3):

x^6 ?

OpenStudy (irish_waffl3):

x^1.5

OpenStudy (irish_waffl3):

not possible

OpenStudy (irish_waffl3):

idk

OpenStudy (welshfella):

No you are DIVIDING so you SUBTRACT the exponents x^3 / x^2 = x^(3-2) = x^1 = x

OpenStudy (irish_waffl3):

oh! when you divide variables, you subtract exponents?

OpenStudy (welshfella):

yes and when yo MULTIPLY variables you ADD exponents

OpenStudy (irish_waffl3):

ok!!! thanks!

OpenStudy (welshfella):

y^2 / y^2 is 1 of course because when you divide something by itse;f anwser is always 1

OpenStudy (irish_waffl3):

correct

OpenStudy (welshfella):

so the second term in the parentheses is -5x

OpenStudy (irish_waffl3):

ok

OpenStudy (irish_waffl3):

so 2x^2y^3+5x?

OpenStudy (welshfella):

so we have 3x^2y^2 ( -2 x^2y^3 - 5x )

OpenStudy (irish_waffl3):

ok

OpenStudy (welshfella):

try and get the third term its + 9x^2y^3 / 3x^2y^2

OpenStudy (welshfella):

9 / 3 = x^2 / x^2 = y^3 / y^2 =

OpenStudy (irish_waffl3):

9/3 is 3

OpenStudy (irish_waffl3):

x^2/x^2 is x

OpenStudy (irish_waffl3):

no its 0

OpenStudy (welshfella):

NO

OpenStudy (welshfella):

its 1

OpenStudy (irish_waffl3):

oh ok

OpenStudy (irish_waffl3):

so 3, 1, and y

OpenStudy (welshfella):

well you needn't put the 1 . its + 3y

OpenStudy (irish_waffl3):

ok

OpenStudy (welshfella):

- thats the last term in the parentheses.

OpenStudy (irish_waffl3):

ok

OpenStudy (irish_waffl3):

thank you! how do I medal?

OpenStudy (welshfella):

so thats it

OpenStudy (welshfella):

press Best Response

OpenStudy (irish_waffl3):

ok, i did! :)

OpenStudy (welshfella):

I like ur name!! lol

OpenStudy (irish_waffl3):

i like yours!

OpenStudy (irish_waffl3):

see ya later fella!

OpenStudy (welshfella):

ok bye

OpenStudy (irish_waffl3):

lol

OpenStudy (irish_waffl3):

bye:)